# OpenLKA: An Open Dataset of Lane Keeping Assist from Recent Car Models under Real-world Driving Conditions

### v1

@@ Abstract
Lane Keeping Assist (LKA) is widely adopted in modern vehicles, yet its real-world performance remains underexplored due to proprietary systems and limited data access. This paper presents OpenLKA, the first open, large-scale dataset for LKA evaluation and improvement. It includes 400 hours of driving data from 50+ production vehicle models, collected through extensive road testing in Tampa, Florida and global contributions from the Comma.ai driving community. The dataset spans a wide range of challenging scenarios, including complex road geometries, degraded lane markings, adverse weather, lighting conditions and surrounding traffic. The dataset is multimodal, comprising: i) full CAN bus streams, decoded using custom reverse-engineered DBC files to extract key LKA events (e.g., system disengagements, lane detection failures); ii) synchronized high-resolution dash-cam video; iii) real-time outputs from Openpilot, providing accurate estimates of road curvature and lane positioning; iv) enhanced scene annotations generated by Vision Language Models, describing lane visibility, pavement quality, weather, lighting, and traffic conditions. By integrating vehicle-internal signals with high-fidelity perception and rich semantic context, OpenLKA provides a comprehensive platform for benchmarking the real-world performance of production LKA systems, identifying safety-critical operational scenarios, and assessing the readiness of current road infrastructure for autonomous driving. The dataset is publicly available at: https://github.com/OpenLKA/OpenLKA.
Index Terms:
Lane Keeping Assist, Autonomous Driving Dataset, CAN–Video Sync Logging, Automatic dataset annotation, Real-World Driving Data
Figure 1: OpenLKA Dataset Overview
@@ I Introduction
To bridge this gap we introduce OpenLKA, an openly available dataset that fuses full CAN messages, high-resolution dash-cam video and lane-detection outputs from production vehicles; the corpus combines a controlled campaign—19 vehicle models driven by three professional drivers for 2 986 LKA-engaged trips (19 380 min) in Tampa, Florida—with 775 community-contributed drives (3 964 min) from 52 additional models that interleave LKA operation and purely human control, thereby yielding in total 61 car models, 95 drivers, 3 761 drives and 400 hours of synchronised CAN–video data, a scale that not only exposes brand- and environment-dependent LKA behaviour but also enables direct human-versus-automation comparisons and rich analyses of driver–LKA interaction in everyday traffic, as summarised in Figure 2.
@@
    A first-of-its-kind open dataset for real-world LKA analysis: OpenLKA provides 400 hours of multimodal driving data from 50+ production vehicles models across various car makes, collected under diverse and challenging real-world conditions, including adverse weather, complex road geometries, and degraded lane markings.
@@
    Reverse-engineering black-box OEM signals for LKA introspection: We develop and apply custom reverse-engineered DBC files to decode internal LKA-related signals from raw CAN bus data—such as system disengagements, lane detection failures, and human driver steering takeover—which are otherwise inaccessible due to the proprietary nature of commercial vehicle systems.
@@
    A unified LKA data collection pipeline and ecosystem: We introduce a scalable data collection framework that captures full CAN bus streams, synchronized dash-cam video, and real-time outputs from Openpilot, enabling scalable estimation of road curvature and lane positioning that would otherwise require costly LIDAR or manual surveys. This pipeline further powers an automated ecosystem that continuously ingests new LKA and human steering data from global Comma.ai community drivers operating in diverse locations and road conditions.
  4. (iv)
@@ II Related Work
Currently, the design and evaluation of ADAS, LKA, and autonomous driving technologies are critically hindered by the lack of comprehensive, multimodal datasets that capture diverse real-world conditions [3, 2]. Existing datasets, such as KITTI [19] and BDD100K [20], primarily focus on sensor and video data but lack crucial CAN bus integration, limiting their applicability in analyzing detailed vehicle dynamics and internal system states critical to understanding LKA behavior. Moreover, many of these datasets are narrowly scoped toward fully autonomous driving scenarios, for example NuScenes  [21], neglecting nuanced human-LKA interactions and failing to represent diverse global driving conditions adequately. This shortcoming restricts researchers from conducting comparative studies between human-driven and automation-driven behaviors, essential for advancing safety and reliability. Additionally, most existing datasets are limited in environmental variability, insufficiently capturing adverse weather, degraded road infrastructure, and complex lighting conditions that are prevalent in real-world driving scenarios. Such limitations impede the development and validation of robust LKA systems capable of handling real-world operational complexities.
Figure 2: Vehicle Distribution in the OpenLKA Dataset
@@ III Experiments, Data Collection and Processing
OpenLKA addresses the critical need for real-world, multimodal, scalable datasets for LKA research. Key design principles include extensibility, openness, multimodal richness and contextual diversity. Unlike existing datasets, OpenLKA integrates synchronized CAN bus, dashcam video, openpilot outputs, and contextual scene annotations, making it uniquely comprehensive.
@@ III-A Experiment Hardware Setup
To evaluate the real-world performance of LKA systems in consumer vehicles [22, 23], we utilized the Comma Three-X [24], an aftermarket self-driving development kit. The Comma Three-X was selected for its integrated high-definition front-view camera and access to the Openpilot ecosystem, which facilitates synchronized data collection. The device connects to the vehicle’s internal Controller Area Network (CAN) bus via the physical On-Board Diagnostics (OBD-II) port, enabling the extraction of high-frequency control signals through reverse engineering, as standard OBD-II protocols do not provide such detailed data.
@@
Figure 3: Comma Three-X dash cam setup for LKA data collection, mounted at the center of the vehicle’s windshield.
@@ III-B Selection of Testing Vehicles and Scenarios
Figure 4: Different of test vehicles across manufacturers.
@@
Figure 5: LKA data collection routes in Tampa, Florida, covering I-275, I-75, and two-lane rural road.
Figure 6: Examples of testing scenarios.
@@ III-C CAN Log Data Acquisition and Reverse Engineering
CAN logs were acquired using the Comma Three-X by accessing the vehicle’s internal CAN bus through the OBD-II port. The logs were recorded in two formats: qlog files (10 Hz, for low-frequency state monitoring) and rlog files (100 Hz, for high-frequency dynamic analysis). We employed open-source DBC files from OpenDBC to decode CAN messages, extracting variables such as clock, LKA/ACC status, pose angles, steering metrics, lane detection, lane offset, object detection, dynamics, and jerk. To ensure accuracy, we manually validated DBC files for Toyota, Kia, and Ford models via signal injection and observation, while Tesla models (e.g., Model 3, Model X) required reverse engineering for precise control signals. Signal availability per vehicle model is detailed in Table II.
Figure 7: Two examples from OpenLKA, the left are the video frames captured by OpenPilot’s dash-cam; the middle are the lane position estimate by the OpenPilot’s end-to-end model; The right are the lane detections results from CAN.
@@
III-D Scalable Lane Position Estimation via Openpilot
We integrate Openpilot, a state-of-the-art open-source advanced driver assistance system, into our dash-cam setup to capture high-fidelity perception outputs alongside front-facing video. Openpilot’s robust vision-based model estimates critical lane-keeping parameters in real time, including the vehicle’s lateral offset from lane markings and the curvature of the road.
These estimations serve as effective proxies for ground truth in our dataset. While not perfectly precise, they offer a practical and scalable alternative to traditional methods such as LiDAR-based lane detection or manual road geometry surveys, which are significantly more costly and labor-intensive. Openpilot thus enables efficient collection of supervisory signals across large-scale, real-world driving data.
Figure 8: In (a) and (b), the left images are the different time driving in similar place captured by Openpilot’s dash-cam, and the middle and right part are both from Openpilot’s output.
By running Openpilot in a passive, observational mode, we preserve the native behavior of the stock LKA system while simultaneously generating consistent, model-based perception outputs. This dual-view setup enables side-by-side comparisons between proprietary systems and a common reference model, supporting fine-grained analysis of lane centering and lateral control performance. An example is shown in Fig. 8, where the vehicle’s lateral deviation from the lane center is estimated and plotted. Manual inspection confirms that the temporal gradient of this deviation closely matches the vehicle’s measured lateral velocity, suggesting strong internal consistency and high precision in Openpilot’s perception outputs. This empirical alignment supports the use of Openpilot as a practical proxy for ground-truth lane position in large-scale observational studies.
III-E VLM-Based Rich Contextual Annotations
@@
Lane Mark. Worn out Partially worn Intact - - - -
Weather Rain Dusty Clear Snow - - -
Lighting Dark Artificial light Normal light - - - -
Traffic Traffic jam Stop-and-go Congested Free flow - - -
Road Cond. Large potholes Small cracks Smooth - - - -
Driving Reckless Major anomaly Minor anomaly Normal - - -
Pedestrian None Few Moderate Crowded - - -
Road Type Two-lane Three-lane Four-lane Multi-lane Roundabout Intersection -
Scenarios None Oncoming vehicle Emergency vehicle Construction Obstacle Animal Accident
Surr. Veh. None Low density Moderate density High density - - -
To provide richer contextual information for detailed analysis of LKA scenarios descriptions, we annotated the OpenLKA video data using VLM, specifically GPT-4o [25]. The video footage was segmented into individual frames, each analyzed together with its immediate preceding and subsequent frames. This method reduces visual obstructions (e.g., windshield wipers, water droplets) and enhances temporal context, leading to more robust and accurate annotations (see Fig. 9).
We developed a domain-specific annotation prompt tailored explicitly for transportation scenarios within the OpenLKA dataset. The prompt utilized System Prompting, Role Prompting and few-shot prompting techniques, instructing the VLM to act as an experienced transportation domain expert [26]. Additionally, the prompt incorporated the Chain-of-Thought (CoT) methodology, promoting systematic, step-by-step reasoning to improve consistency and reduce annotation hallucinations [27]. Few-shot examples were also included to clearly illustrate the intended annotation process. The VLM classified each frame into predefined environmental categories (as shown in Table III). Annotation accuracy was validated by manually labeling a representative subset of 2,000 images, selected via stratified random sampling across all categories. Comparing manual annotations with GPT-4o’s outputs revealed an accuracy exceeding 85%. Notably, the VLM exhibited a conservative labeling tendency, frequently identifying scenarios as more severe than manual annotations, a beneficial characteristic for safety-critical transportation system design.
@@ IV Dataset Overview
CAN & Openpilot System Status Activation states for Openpilot, adaptive cruise control, and lane-keeping systems.
Vehicle Dynamics Vehicle motion data including speed, acceleration, steering, and body orientation.
Perception Outputs Environmental awareness data such as lane visibility, lane departure risks, and lead vehicle detection.
Planning Outputs Path planning data including desired curvature and predicted trajectories.
@@
The OpenLKA dataset was developed to address the need for a comprehensive, multimodal resource tailored for LKA system research, offering a diverse and parameter-rich foundation for analyzing real-world LKA performance and human-LKA interactions. Spanning 389.07 hours across 1,166 unique driving routes, the dataset encompasses data from 50+ vehicle models and nearly 100 drivers, organized in a structured car model, dongle id, route id format to distinguish between models, drivers, and trips. In the first, self-collected LKA data, was gathered starting in May 2024 by three drivers using 19 market-leading vehicle models rented from Tampa International Airport, covering a wide range of brands prevalent in the U.S. market, resulting in 323 hours (83% of the total duration) of driving data across Florida’s diverse weather, lighting, road, and traffic conditions; the second component, sourced from the Comma autonomous driving community, comprises 66.1 hours (17%) of human driving data from 92 drivers across 50+ vehicle models worldwide, collected as of April 2025, providing a valuable baseline for comparing human driving styles with LKA performance and exploring human-LKA interaction dynamics, as illustrated by the global distribution in Figure 10 [24]. Additionally, this part of the dataset covers more than 30 states, provinces, and territories in over 80 cities around the world. This greatly enriches the dataset with driving scenarios such as snow, dust storms, and frosty roads, as well as driving scenarios in mountainous areas, dense cities, and deserts.
Figure 10: Global Distribution of Community Driving Dataset
The dataset’s diversity is evident in its composition, the most represented model, HYUNDAI IONIQ 5 2022, contributes 14.61% of the total duration (56.85 hours). The dataset includes synchronized multimodal streams, such as videos at multiple resolutions, vehicle control data via CAN, Openpilot’s end-to-end model outputs, and VLM-annotated environmental factors, with a detailed breakdown provided in Table IV. Privacy is ensured by blurring faces and license plates in all videos to protect driver identities.
Figure 11: LKA performance overview under normal conditions
Table V compares OpenLKA with existing autonomous driving datasets, underscoring its unique strengths: comprehensive environmental annotations covering weather, visibility, traffic, and infrastructure; synchronized video and CAN data for precise temporal alignment; rich CAN signals paired with Openpilot’s baseline outputs; and VLM-augmented labels that enhance contextual understanding, positioning OpenLKA as a uniquely diverse and parameter-rich resource for advancing LKA system development and dataset-driven research in autonomous driving.
@@
V Dataset Applications

### v2

# OpenLKA: An Open Dataset of Lane Keeping Assist from Recent Production Vehicles under Real-World Driving Conditions

@@ Abstract
Lane Keeping Assist (LKA) is widely adopted in modern production vehicles, yet its real-world performance remains opaque due to the system’s proprietary control stack, preventing researchers from diagnosing or improving this technology. To fill the gap, this paper presents OpenLKA, the first open, large-scale dataset for LKA evaluation and improvement. It includes 389.1 hours of LKA-steered data from 62 production vehicle models, collected through extensive road testing in Tampa, Florida and global contributors from the open-source community. The dataset spans a wide range of challenging scenarios, including degraded lane markings, complex road geometries, adverse weather, lighting conditions and various surrounding traffic. The dataset is multimodal, comprising: i) decoded vehicle internal messages including key LKA signals (e.g., system disengagements, lane detection failures); ii) synchronized high-resolution videos from a mounted dash camera; and iii) rich scene annotations generated by a vision language model, covering lane visibility, pavement quality, weather, lighting, and traffic conditions etc. Collectively, OpenLKA provides a comprehensive platform for benchmarking the real-world performance of production LKA systems, identifying safety-critical operational scenarios, and assessing the readiness of current road infrastructure for autonomous driving. The dataset is publicly available at: https://github.com/OpenLKA.
Figure 1: OpenLKA dataset overview [^2] Logos © respective manufacturers (media-kit downloads, May 2025); icons © The Noun Project, $C$ BY 3.0..
@@ I Introduction
To bridge this gap we introduce OpenLKA; see Fig.2;, an openly available dataset that fuses full Controller Area Network (CAN) messages, high-resolution dash-cam video and lane-detection outputs from production vehicles; the corpus combines a controlled campaign—19 vehicle models driven by three professional drivers for 2,986 LKA-engaged trips (19,380 min) in Tampa, Florida—with 775 community-contributed drives (3,964 min) from 52 models that interleave LKA operation and purely human control, thereby yielding in total 62 car models, 52 drivers, 3,761 drives and 389.1 hours of synchronised CAN–video data, a scale that not only exposes brand- and environment-dependent LKA behaviour but also enables direct human-versus-automation comparisons and rich analyses of driver–LKA interaction in everyday traffic, as summarised in Figure 2.
@@
    A first-of-its-kind open dataset for real-world LKA analysis: OpenLKA provides 389.1 hours of multimodal driving data from 62 production vehicle models across various car makes, collected under diverse and challenging real-world conditions, including adverse weather, complex road geometries, and degraded lane markings.
@@
    A unified LKA data collection pipeline and ecosystem: We introduce a scalable data collection framework that captures full CAN bus streams and synchronized dash-cam video, together with real-time lane detection results from a parallel vision model from Openpilot [13], which enables scalable estimation of road curvature and vehicle in-lane position that would otherwise require costly LIDAR or manual surveys. This pipeline further powers an ecosystem that continuously ingests new LKA and human steering data from global community drivers operating in diverse locations and road conditions.
@@ II Related Work
Currently, the design and evaluation of ADAS, LKA, and autonomous driving technologies are critically hindered by the lack of comprehensive, multimodal datasets that capture diverse real-world conditions [3, 2]. Existing datasets, such as KITTI [20] and BDD100K [21], primarily focus on sensor and video data but lack crucial CAN bus integration, limiting their applicability in analyzing detailed vehicle dynamics and internal system states critical to understanding LKA behavior. Moreover, many of these datasets are narrowly scoped toward fully autonomous driving scenarios in single specific testing vehicle, for example NuScenes [22], neglecting nuanced human-LKA interactions and failing to represent diverse global driving conditions adequately. This shortcoming restricts researchers from conducting comparative studies between human-driven and automation-driven behaviors, essential for advancing safety and reliability. Additionally, most existing datasets are limited in environmental variability, insufficiently capturing adverse weather, degraded road infrastructure, and complex lighting conditions that are prevalent in real-world driving scenarios. Such limitations impede the development and validation of robust LKA systems capable of handling real-world operational complexities.
Figure 2: Distribution of hours by car models in OpenLKA
@@ III Experiments, Data Collection and Processing
OpenLKA addresses the critical need for real-world, multimodal, scalable datasets for LKA research. Key design principles include extensibility, openness, multimodal richness and contextual diversity. Unlike existing datasets, OpenLKA integrates synchronized CAN bus, dashcam video, OP outputs, and contextual scene annotations, making it uniquely comprehensive.
@@ III-A Experiment Hardware Setup
Figure 3: Experiment setup for LKA data collection and testing fleet using recent rental cars from several car makes.
To evaluate the real-world performance of LKA systems in production vehicles [23, 24], we utilized the Comma Three-X [25], an aftermarket self-driving development kit. The Comma Three-X was selected for its integrated high-definition front-view camera and access to the OP ecosystem, which facilitates synchronized data collection. The device connects to the vehicle’s internal Controller Area Network (CAN) bus via the physical On-Board Diagnostics (OBD-II) port, enabling the extraction of high-frequency control signals through reverse engineering, as standard OBD-II protocols do not provide such detailed data.
@@ III-B Selection of Testing Vehicles and Scenarios
Figure 4: LKA data collection sites in Tampa, Florida, covering I-275, I-75, and two-lane, two-way rural road.
Figure 5: Examples of testing scenarios in OpenLKA.
@@ III-C CAN Log Data Acquisition and Reverse Engineering
CAN logs were acquired using the Comma Three‑X by accessing the vehicle’s internal CAN bus through the OBD‑II port. The logs were recorded in two formats: qlog files (10 Hz, for low‑frequency state monitoring) and rlog files (100 Hz, for high‑frequency dynamic analysis). We initially employed open‑source DBC files from OpenDBC to decode CAN messages—extracting variables such as clock, LKA/ACC status, pose angles, steering metrics, lane detection, lane offset, object detection, dynamics, and jerk. The right part of Fig. 6 shows an example that we reversed engineer the laneline detection results and LKA status from Hyundai IONIQ5. The availability of signals per vehicle model is detailed in Table II. However, OpenDBC files only natively support a subset of vehicle makes (e.g., Ford, Toyota Camry, Corolla, Kia, Chevrolet). To extend support to remaining vehicle models, we performed additional work: (1) curated and integrated richer DBC definitions by retrieving files shared in public forums and inferring message structures from partially documented CAN signal information; (2) reverse‑engineered DBCs by adapting files from similar or earlier‑model years of the same make; and (3) applied a systematic process of coarse‑to‑fine message localization—grouping signals by message ID header, correlating with known vehicle behavior domains, and validating candidate signals via controlled vehicle observations. All resulting DBC files have been openly released in our GitHub repository to support future analysis and transparency.
Figure 6: CAN and video data sample collected in heavy rain: (a) the front camera view. (b) LKA-related CAN data in 60s, including lane deviation (blue), laneline detection result (green, orange; 1 solid, 2 faded, 3 departure), laneline detection confidences by OP (magenta, red), and LKA engagement: active (blue shading) vs. disengaged (grey).
@@
III-D Scalable and Robust Laneline Estimation via OP
We integrate OP, a state‑of‑the‑art, open‑source advanced driver‑assistance system, into our dash‑cam platform to record high‑fidelity perception outputs that are time‑aligned with the forward‑facing video. OP’s vision model continuously estimates key lane‑keeping variables, including the vehicle’s lateral offset to each lane line and the instantaneous road curvature.
Although not perfectly exact, these estimates provide a quite accurate, practical, scalable proxy for ground truth, avoiding the cost and labor of LiDAR‑based mapping or manual road‑geometry surveys. OP thus enables efficient collection of supervisory signals across large‑scale, real‑world driving data. Figure 7 plots the OP‑estimated distances to the left and right lane lines; their sum approximates the roadway width, which remains stable throughout the drive. Thanks to OP’s end‑to‑end architecture, its lane‑detection confidence is consistently higher than that of the stock system, as shown in Fig. 6.
Running OP in a passive, observational mode preserves the native behavior of the production LKA while simultaneously generating a consistent, model‑based perception stream. This dual‑view setup enables side‑by‑side comparisons between proprietary controllers and a common reference model, supporting fine‑grained analyses of lane centering and lateral control. The middle panels of Fig. 6 illustrate an example: OP directly yields the lateral distances from the vehicle center to the left and right lane boundaries (positive to the right), from which we compute the lane‑center deviation. Manual inspection confirms that the temporal derivative of this deviation closely matches the measured lateral velocity, indicating strong internal consistency and high precision in OP’s perception outputs. This empirical agreement justifies using OP estimates as a practical surrogate for ground‑truth lane position in large‑scale observational studies.
Figure 7: Top: OP‑estimated lateral distances from the vehicle centre to the left (blue) and right (orange) lane lines. Bottom: Road width (green) obtained by summing the two distances; the dashed red line marks the nominal 3.35 m lane width. The width remains nearly constant, suggesting the robustness and accuracy of OP’s lane detection algorithms.
III-E VLM-Based Scene Annotations
@@
Lane Mark. Worn out Partially worn Intact
Weather Rain Dusty Clear Snow
Lighting Night Night with light Daylight
Traffic Traffic jam Stop–and–go Congested Free flow
Road Cond. Large potholes Small cracks Smooth
Driving Reckless Aggressive Normal Timid
Pedestrian None Few Moderate Crowded
Road Type Two–Lanea Three–Laneb Four–Lanec Multilaned Laneline Transition Intersection
Scenarios None Oncoming vehicle Emergency vehicle Construction Obstacle Accident Others
Surr. Veh. None Low density Moderate density High density
  • a
    2U: Two–lane, two–way undivided road.
  • b
    3T: Three–lane, two–way road with center two–way left–turn lane (TWLTL).
  • c
    4U: Four–lane, two–way undivided arterial.
  • d
[…]
@@
To provide richer contextual information for detailed analysis of LKA scenarios descriptions, we annotated the OpenLKA video data using VLM [26], specifically GPT-4o. The video footage was segmented into individual frames, each analyzed together with its immediate preceding and subsequent frames. This method reduces visual obstructions (e.g., windshield wipers, water droplets) and enhances temporal context, leading to more robust and accurate annotations (see Fig. 8).
We developed a domain-specific annotation prompt tailored explicitly for transportation scenarios within the OpenLKA dataset. The prompt utilized System Prompting, Role Prompting and few-shot prompting techniques, instructing the VLM to act as an experienced transportation domain expert [27]. Additionally, the prompt incorporated the Chain-of-Thought (CoT) methodology[28], promoting systematic, step-by-step reasoning to improve consistency and reduce annotation hallucinations [28]. Few-shot examples were also included to clearly illustrate the intended annotation process. The VLM classified each frame into predefined environmental categories (as shown in Table III). Annotation accuracy was validated by manually labeling a representative subset of 2,000 images, selected via stratified random sampling across all categories. Comparing manual annotations with GPT-4o’s outputs revealed an overall accuracy exceeding 94% and each category’s accuracy exceeding 85%, shown in Fig. 9. Notably, the VLM exhibited a conservative labeling tendency, frequently identifying scenarios as more severe than manual annotations, a beneficial characteristic for safety-critical transportation system design.
Figure 9: Comparision of human’s label and VLM’s annotation on sampled dataset.
@@ IV Dataset Overview
Decoded CAN LKA Status LKA System engage status, Lane depature warning output, OEM detection outputs
Vehicle Dynamics Motion data including speed, acceleration, steering, and body orientation.
Vision Detection Model(OP) Outputs Perception Outputs Environmental awareness data, e.g. lane visibility, lane departure risks, and lead vehicle detection.
Planning Outputs Path planning, e.g. desired curvature and predicted trajectories.
@@
The OpenLKA dataset was developed to address the need for a comprehensive, multimodal resource tailored for LKA system research, offering a diverse and parameter-rich foundation for analyzing real-world LKA performance and human-LKA interactions. Spanning 389.1 hours across 1,166 unique driving routes, the dataset encompasses data from 62 vehicle models and 52 drivers, organized in a structured car model, dongle id, route id format to distinguish between models, drivers, and trips. In the first, self-collected LKA data, was gathered starting in May 2024 by three drivers using 19 market-leading vehicle models rented from Tampa International Airport, covering a wide range of brands prevalent in the U.S. market, resulting in 323 hours (83% of the total duration) of driving data across Florida’s diverse weather, lighting, road, and traffic conditions; the second component, sourced from the Comma autonomous driving community, comprises 66.1 hours (17%) of human driving data from 49 drivers across 50+ vehicle models worldwide, collected as of April 2025, providing a valuable baseline for comparing human driving styles with LKA performance and exploring human-LKA interaction dynamics, as illustrated by the global distribution in Figure 10 [25]. Additionally, this part of the dataset covers more than 30 states, provinces, and territories in over 80 cities around the world. This greatly enriches the dataset with driving scenarios such as snow, dust storms, and frosty roads, as well as driving scenarios in mountainous areas, dense cities, and deserts.
Figure 10: World-wide open driving logs from community
The dataset’s diversity is evident in its composition, the most represented model, HYUNDAI IONIQ 5 2022, contributes 14.61% of the total duration (56.85 hours). The dataset includes synchronized multimodal streams, such as videos at multiple resolutions, vehicle control data via CAN, OP’s end-to-end model outputs, and VLM-annotated environmental factors, with a detailed breakdown provided in Table IV. Privacy is ensured by blurring faces and license plates in all videos to protect driver identities.
Table V compares OpenLKA with existing autonomous driving datasets, underscoring its unique strengths: comprehensive environmental annotations covering weather, visibility, traffic, and infrastructure; synchronized video and CAN data for precise temporal alignment; rich CAN signals paired with OP’s baseline outputs; and VLM-augmented labels that enhance contextual understanding, positioning OpenLKA as a uniquely diverse and parameter-rich resource for advancing LKA system development and dataset-driven research in autonomous driving.
@@
By uniting production‑vehicle data collection with tightly synchronized video, CAN, OP outputs, and VLM scene labels, OpenLKA delivers a cost‑efficient yet richly annotated mirror of everyday lane‑keeping. This breadth of signals and operating conditions not only establishes a realistic benchmark for today’s LKA systems, but also unlocks a wide spectrum of downstream studies—from algorithm refinement to infrastructure evaluation and human–machine interaction analysis. The following Discussion illustrates these opportunities and charts concrete research directions enabled by the dataset.
V Discussion: Dataset Applications